AB INITIO REACTION PATHWAYS FOR β− HYDROGEN ELIMINATION FROM SMALL RADICALS

L. MENDOZA, E. MIKESELL, L. NGUYEN and \underline{P. E. FLEMING}, Department of Chemistry, California State University at Sacramento, Sacramento, CA 95819.

Reaction pathways have been calculated for the elimination of a β− hydrogen from several small radicals. Structural and thermodynamic information is predicted by the calculations. Calculated structural changes are consistent with very simple models of chemical bonding involving a change in the orbital hybridization of the large fragment atom involved in the bond that is dissociated. Calculations at the MP2/6-311g* level give thermodynamic values which are in modest agreement with experiment.